The photoelectric threshold frequency of platinum is 1.54 ✕ 1015 Hz. What is the minimum energy required to remove an electron from this element?
Answer #1

here,

The photoelectric threshold frequency of platinum , f = 1.54 * 10^15 Hz

the minimum energy required to remove an electron from this element , U = h * f

U = 6.634 * 10^-34 * 1.54 * 10^15 J

U = 1.02 * 10^-18 J

the minimum energy required to remove an electron from this element is 1.02 * 10^-18 J